OpenStudy (anonymous):

Find S11 for 3 + 1 + (-1) + (-3) +…

OpenStudy (anonymous):

so the 1st term A = 3, Common difference D = 1-3 = -2, n = 11 So 1st you have to find the 11th term. Tn = A + (n-1)D so find T11 1st .

OpenStudy (anonymous):

11=3+(11-2)1-3

OpenStudy (anonymous):

oh LOL. T11 means 11th term. T11 = 3 + (11-1).-2 a = 3, n = 11, d = -2.

OpenStudy (anonymous):

Answers are -77 -75 -73 -71

OpenStudy (anonymous):

Im not getting none of these answers

OpenStudy (anonymous):

Ok. T11 means 11th term which is 3 + (11-1)-2 => 3 -20 = -17. so a = 3, d = -2, and T11 = -17 If we have N Terms, then sum of N terms => N(a+Nth term)/2 Similarly u can find S11 = 11(3-17)/2, and see the ans.

OpenStudy (anonymous):

-77=A :)